Date Shifter Run Information Detector Report Defects Comment
| | | | | | |

01-06-2016 laurent chevalier Run 300600
Project Tag: data16_13TeV
Run Start: 2016-05-31 18:02 CEST
Run End: 2016-06-01 08:09 CEST
# lumi blocks: 870
Events: 15389940
Stable beams: Yes
Online ATLAS Ready luminosity: 59.965 pb-1
CSC
CSC EA no hits in A07, layer 1, channels 85-97 link
CSC EC 1 & 3 known problems link
RPC A
Side A RPC always noisy: BOF 4 &BOG 3 (phi sector 11 & 12) link
RPC C
Missing hits on: BM2, BMH2C02, BML1C05HR, BMS2C08HV, BMH2C08HV,BML1C13& BMS2C02 on the RO side link
TGC side A
wire occupancy L2_E3, same as before
TGC side C
noise around C10 & C02 link
Tracking
[eta~0,phi~0](BMS2C02) + [eta~-0.5,phi~2.7](?) + [eta~0.5,-0.5<phi<0.5] (?) see link
Tolerable: MS-CSC-EC-DISCONNECTED2

01-06-2016 laurent chevalier Run 300571
Project Tag: data16_13TeV
Run Start: 2016-05-31 06:09 CEST
Run End: 2016-05-31 16:10 CEST
# lumi blocks: 627
Events: 33694798
Stable beams: Yes
Online ATLAS Ready luminosity: 164.550 pb-1
CSC
CSC EA no hits in A07, layer 1, channels 85-97 link
CSC EC 1 & 3 known problems link
RPC A
Side A RPC always noisy: BOF 4 &BOG 3 (phi sector 11 & 12) link
RPC C
Missing hits on: BM2, BMH2C02, BML1C05HR, BMS2C08HV, BMH2C08HV,BML1C13& BMS2C02 on the RO side link
TGC side A
wire occupancy L2_E3, same as before
TGC side C
noise around C10 & C02 link
Tracking
[eta~0,phi~0](BMS2C02) + [eta~-0.5,phi~2.7](?) + [eta~0.5,-0.5<phi<0.5] (?) see link
Tolerable: MS-CSC-EC-DISCONNECTED2

30-05-2016 laurent chevalier Run 300540
Project Tag: data16_13TeV
Run Start: 2016-05-30 12:37 CEST
Run End: 2016-05-31 04:52 CEST
# lumi blocks: 1004
Events: 40854328
Stable beams: Yes
Online ATLAS Ready luminosity: 212.867 pb-1
Remarks: no infos for LB: ~350, ~400, ~780
Same problem as before:
MS-CSC-EC-DISCONNECTED2
CSC
CSC EA no hits in A07, layer 1, channels 85-97 link
CSC EC 1 & 3 known problems link
RPC
Missing hits on: BMH2C02& BMS2C02 on the RO side link see also expert plot
RPC always noisy: BOF 4 &BOG 3 (phi sector 11 & 12) link
TGC side A
wire occupancy L2_E3, same as before
TGC side C
noise around C10 & C02 link
strips and wires occupancy same as before (run# 300415)
Tracking
[eta~0,phi~0](BMS2C02) + [eta~-0.5,phi~2.7](?) + [eta~0.5,-0.5<phi<0.5] (?) see link
 

30-05-2016 laurent chevalier Run 300487
Project Tag: data16_13TeV
Run Start: 2016-05-29 18:57 CEST
Run End: 2016-05-30 07:38 CEST
# lumi blocks: 780
Events: 32538443
Stable beams: Yes
Online ATLAS Ready luminosity: 147.527 pb-1
some problems
MS-CSC-EC-DISCONNECTED2
CSC 1 & 3 known problems
CSC EA no hits in A07, layer 1, channels 85-97 link
*TGCsideA*
noisy in A02EIFI see link
and wire occupancy L2_E3, same as before
*TGCsideC*
noise around C10 & C02 link
strips and wires occupancy same as before (run# 300415)
 

29-05-2016 laurent chevalier Run 300418
Project Tag: data16_13TeV
Run Start: 2016-05-28 22:11 CEST
Run End: 2016-05-29 03:16 CEST
# lumi blocks: 316
Events: 2691040
Stable beams: Yes
Online ATLAS Ready luminosity: 6.903 pb-1
some problems
MS-CSC-EC-DISCONNECTED2
RPC: BOF 4 &BOG 3 (phi sector 11 & 12 , feet chambers) & BME1 missing hits link
TGC: noise around C10 & C02 link
Tracking, fewer pixel hits per combined muon link
   

29-05-2016 laurent chevalier Run 300415
Project Tag: data16_13TeV
Run Start: 2016-05-28 11:50 CEST
Run End: 2016-05-28 21:55 CEST
# lumi blocks: 645
Events: 25973953
Stable beams: Yes
Online ATLAS Ready luminosity: 107.523 pb-1
some problems
MS-CSC-EC-DISCONNECTED2
CSC EA no hits in A07, layer 1, channels 85-95 link
RPC BA Low pT phi-panel hot spot in BOF 4 &BOG 3 (phi sector 11 & 12 , feet chambers) & BME1 missing hits link
TGC noisy in C10 Layer 4 E 3,4,5 & C02link Muon Tracking
Number of Zs v. LB Jumps around LB 200 link
number of Muons/event v. LB is constant link
Increase in event rate due to a change in lumi by LHC - see trigger shifter elog
CSC C01 has fewer segments per chamber compared to other CSC chambers, and this is expected because of the HV problems link

Tolerable: 2 layers disconnected in one sector.
Coverage loss < 10% two layers in sector C01 have no hits from LB 200-END
 

28-05-2016 laurent chevalier Run 300345
Project Tag: data16_13TeV
Run Start: 2016-05-27 20:13 CEST
Run End: 2016-05-28 00:43 CEST
# lumi blocks: 282
Events: 2514669
Stable beams: Yes
Online ATLAS Ready luminosity: 6.851 pb-1
everything seems ok, small remarks
RPC problems: see link
BME1C13
BMS2C02 - RO side
BMH2C0 - RO side
BML4C07 - HV Side
BML7C11
   

28-05-2016 laurent chevalier Run 300312
Project Tag: data16_cos
Run Start: 2016-05-27 16:55 CEST
Run End: 2016-05-27 17:01 CEST
# lumi blocks: 7
Events: 253511
Stable beams: No
Online ATLAS Ready luminosity: 0.000 pb-1
Only cosmic-Calo
     

28-05-2016 laurent chevalier Run 300287
Project Tag: data16_13TeV
Run Start: 2016-05-26 23:34 CEST
Run End: 2016-05-27 05:54 CEST
# lumi blocks: 399
Events: 3286803
Stable beams: Yes
Online ATLAS Ready luminosity: 9.756 pb-1
everything seems ok, small remarks:
BOA6 noisy, see link
BOA7,8 noisy, see link
BOC6 sector 4 , see link and link
EEA2 (5) off?, see link
EEC2 (5) off? link
EIL3 sector C link
Muons & Jpsi number of segments in EIS & BEE per chamber, ex: link
     

27-05-2016 laurent chevalier Run 300279
Project Tag: data16_13TeV
Run Start: 2016-05-26 23:34 CEST
Run End: 2016-05-27 05:54 CEST
# lumi blocks: 399
Events: 3286803
Stable beams: Yes
Online ATLAS Ready luminosity: 9.756 pb-1
everything seems ok, small remarks:
CSC A07, layer 1 channels 85-95 have no hits link
The number of TRT hits per combined muons was flagged red due to a TRT problem link
   

27-05-2016 laurent chevalier Run 300211
Project Tag: data16_cos
Run Start: 2016-05-25 21:20 CEST
Run End: 2016-05-26 08:49 CEST
# lumi blocks: 687
Events: 5875635
Stable beams: No
everything seems ok, small remarks:
missing RPC A & C-side link
   

27-05-2016 laurent chevalier Run 300204
Project Tag: data16_cos
Run Start: 2016-05-25 19:50 CEST
Run End: 2016-05-25 20:40 CEST
# lumi blocks: 50
Events: 384265
Stable beams: No
everything seems ok, small remarks:
CSC ???? nosy see:link
L1Mu Red: probably too long cosmic run seems ok (sector5 and 13 only)
RPC BOG2C12? noisy see:link
RPC BMF1C12? noisy see:link
   

-- LaurentChevalier - 2016-05-27

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2016-06-01 - LaurentChevalier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback